when fritz drives to work his trip takes 40 minutes, but when he takes the train it takes 30 minutes. find the distance fritz travels to work if the train travels an average of 15 mph faster than his driving.
Answer by checkley71(8403)

D=RT
SEEING AS THE DISTANCES ARE EQUAL THEN
R(40/60)=(R+15)(30/60)
40R/60=(30R+450)/60 HERE WE CAN ELIMINATE THE DENOMINATORS THUS WE HAVE LEFT:
40R-30R=450
10R=450
R=450/10
R=45 MPH FOR THE DRIVE WHILE THE TRAIN'S RATE IS:
45+15=60 MPH FOR THE TRAIN.
PROOF
45(40/60)=60(30/60)
45*40=60*30
1800=1800